A poll is taken in which 342 out of 500 randomly selected voters indicated their preference for a certain candidate.

(a) Find a 98% confidence interval for p.

≤ p ≤

(b) Find the margin of error for this 98% confidence interval for p.
Answer #1

solution:-

given that n = 500 , x = 342

proportion p = x/n = 342/500 = 0.684

(a) the value of 98% confidence from z table is 2.33

confidence interval formula

=> p +/- z* sqrt(p*(1-p)/n)

=> 0.684 +/- 2.33 * sqrt(0.684*(1-0.684)/500)

=> 0.636 ≤ p ≤ 0.732 rounded to three decimals


(b) margin of error formula

=> z* sqrt(p*(1-p)/n)

=> 2.33 * sqrt(0.684*(1-0.684)/500)

=> 0.048 rounded to three decimals